# Reservoir Computing as a Language Model

### v2

@@ Abstract
Large Language Models (LLM) have dominated the science and media landscape duo to their impressive performance on processing large chunks of data and produce human-like levels of text. Nevertheless, their huge energy demand and slow processing still a bottleneck for further increasing quality while also making the models accessible to everyone. To solve this bottleneck, we will investigate how reservoir computing performs on natural text processing, which could enable fast and energy efficient hardware implementations. Studies investigating the use of reservoir computing as a language model remain sparse. In this paper, we compare three distinct approaches for character-level language modeling, two different reservoir computing approaches, where only an output layer is trainable, and the well-known transformer-based architectures, which fully learn an attention-based sequence representation. We explore the performance, computational cost and prediction accuracy for both paradigms by equally varying the number of trainable parameters for all models. Using a consistent pipeline for all three approaches, we demonstrate that transformers excel in prediction quality, whereas reservoir computers remain highly efficient reducing the training and inference speed. Furthermore, we investigate two types of reservoir computing: a traditional reservoir with a static linear readout, and an attention-enhanced reservoir that dynamically adapts its output weights via an attention mechanism. Our findings underline how these paradigms scale and offer guidelines to balance resource constraints with performance.
@@ I Introduction
Modern sequence modeling tasks, such as language modeling and machine translation, have been dominated by attention-based architectures-most prominently the transformer family [1, 2, 3, 4, 5, 6]- which excel at capturing long-range dependencies by learning contextual representations through self-attention layers and feed-forward networks. However, transformer training and inference incur substantial computational and energy costs, often requiring specialized hardware and limiting accessibility with tighter resource budgets.
Reservoir computing (RC) offers an alternative. In RC, a large, fixed, recurrent “reservoir” projects inputs into a high-dimensional state space, and only a lightweight readout layer is trained. This design dramatically reduces training time and energy consumption and can be implemented efficiently in software or on analog hardware substrates. Recent work has extended classical RC with neural programming techniques to shape reservoir dynamics and introduced attention mechanisms at the readout stage to adaptively weight reservoir states [7]. RC has been shown to excel in time-series prediction [8, 9, 10], capitalizing on the mentioned large, fixed random reservoir for capturing nonlinear dynamics with minimal training overhead [11, 12, 13, 14]. The charm of reservoir computing lies in its versatile implementation via a wide range of physical substrates [15], including quantum systems as potential RC candidates [16, 17, 18].
Recently, more expressive control over reservoir behavior has been introduced via neural programming paradigms, further expanding their utility in symbolic computation and structured sequence modeling [19]. While prior work has explored the use of reservoir computing for text classification tasks [20] and even demonstrated hardware-oriented implementations for language learning [21], the use of reservoir computing explicitly as a language model—i.e., for generative next-token prediction—remains largely unexplored. Reservoir computing has an advantage of simplicity, since only a small readout layer must be trained, while transformers require end-to-end training of embeddings, multi-head self-attention layers, and feed-forward networks. However, transformers can achieve state-of-the-art performance in language tasks by scaling to millions (or billions) of parameters [4].
In this study, we present a unified framework, applying both approaches to character-level sequence prediction on a small corpus of Shakespeare text inspired by recent minimalist transformer implementations such as Karpathy’s NanoGPT [22], which demonstrate how compact models can still achieve strong performance on character-level language modeling tasks. We vary reservoir size in the classic RC for a linear readout and RC size and attention layer size for an attention-enhanced readout [7], while varying hidden dimension and number of layers in transformers, obtaining multiple configurations spanning from thousands up to hundred thousand of trainable parameters. By measuring cross-entropy loss across these diverse setups, we expose fundamental trade-offs in resource usage and generalization performance.
@@
Figure 1: Diagram illustrating the reservoir computing model with vector embedding, reservoir processing (Att-Enhanced Reservoir, Transformer), output layer, and predicted probability vector for the next letter. The vector embedding is trained for the transformer case, while being randomly initialized for the reservoirs. The predicted probabilities (e.g., $p_{A}$ for letter a) are used to compute the cross entropy loss $H(y,\hat{y})=-\sum_{i}y_{i}\log\hat{y}_{i}$, which drives the error feedback.
II Task Formulation and Dataset
We tackle a character-level next-character prediction task. Let $\mathbf{X}=\{x_{1},\dots,x_{T}\}$ represent a fixed-length input sequence of characters, and let $y_{T+1}$ be the next character to predict. We treat the problem as a multi-class classification over the vocabulary $\mathcal{V}=59$. The dataset is a 9-million-character corpus of Shakespeare text inspired by recent minimalist transformer implementations such as Karpathy’s NanoGPT [22], which we split into 666 shards: 555 shards for training and 111 shard for testing. Each shard is processed at the character level, forming sequences of length 323232 for the input, plus the next character as the target. Before training the text corpus is preprocessed by lowercasing all letters to achieve a smaller vocabulary $\mathcal{V}=59$. Training optimizes a standard cross-entropy loss over the predicted logits versus the correct next character. We track training and test losses per shard and epoch.
III Background and Model Architectures
III.1 Traditional Reservoir Computing
@@
The echo state implementation of a classic reservoir computer consists of three main components: an input-to-reservoir weight matrix $\mathbf{W}_{\mathrm{in}}\in\mathbb{R}^{N\times d}$, a fixed recurrent reservoir matrix $\mathbf{W}_{\mathrm{res}}\in\mathbb{R}^{N\times N}$, and a readout layer $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{\mathcal{V}\times N}$ [11]. Here $d$ is the dimension of the chosen input-embedding vector for the input letter, $N$ the number of nodes in the reservoir, and $\mathcal{V}$ the size of the vocabulary. The reservoir evolves its hidden states $\mathbf{r}_{t}\in\mathbb{R}^{N}$ as
$\mathbf{r}_{t}=\tanh(\mathbf{r}_{t-1}\,\mathbf{W}_{\mathrm{res}}+\mathbf{x}_{t}\,\mathbf{W}_{\mathrm{in}}).$
@@
Here, $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is an embedding vector of the current input character. In the simplest (traditional) reservoir setting, only the readout layer $\mathbf{W}_{\mathrm{out}}$ is trained:
$\bar{\mathbf{y}}_{t}=\mathbf{r}_{t}\,\mathbf{W}_{\mathrm{out}}.$
@@
where $\mathbf{r}_{t}$ is the reservoir state at time $t$. Typically, $\mathbf{W}_{\mathrm{out}}$ is learned using ridge regression, which minimizes the mean squared error with an added $\ell_{2}$ regularization term to prevent overfitting and improve generalization. However, in our experiments, we employ a cross-entropy loss function, necessitating gradient-based optimization methods for training $\mathbf{W}_{\mathrm{out}}$. Moreover, ridge regression can become computationally intensive as the number of training samples and reservoir size increase, due to the inversion of large matrices. By using gradient-based methods, we circumvent these scalability issues.
@@
Table 1: Trainable Parameter Counts for the classic reservoir, AERC, and Transformer Architectures
@@
Transformer Hidden Size $d_{h}=64$, Heads $h=4$, Layers $L=4$ 150671506715067
Hidden Size $d_{h}=72$, Heads $h=8$, Layers $L=8$ 302993029930299
Hidden Size $d_{h}=128$, Heads $h=8$, Layers $L=8$ 450834508345083
Hidden Size $d_{h}=356$, Heads $h=8$, Layers $L=8$ 105275105275105275
Hidden Size $d_{h}=256$, Heads $h=16$, Layers $L=16$ 155803155803155803
Reservoir Reservoir Size $N=250$ 148091480914809
Reservoir Size $N=500$ 295592955929559
Reservoir Size $N=750$ 443094430944309
Reservoir Size $N=1750$ 103309103309103309
Reservoir Size $N=2600$ 153459153459153459
Att Reservoir Reservoir Size $N=75$, Hidden Size $H=13$ 154641546415464
Reservoir Size $N=75$, Hidden Size $H=19$ 311243112431124
Reservoir Size $N=100$, Hidden Size $H=20$ 452594525945259
Reservoir Size $N=150$, Hidden Size $H=25$ 102809102809102809
Reservoir Size $N=160$, Hidden Size $H=30$ 155459155459155459
III.2 Attention-Enhanced Reservoir Computing
@@
$\mathbf{W}_{\text{att},l}=F(\mathbf{W}_{\text{net}},\mathbf{r}_{l}).$
@@
$\mathbf{r_{o}}_{l}=\mathbf{W}_{\text{att},l}\mathbf{r}_{l}.$
@@
$\bar{\mathbf{y}}_{l}=\mathbf{W}_{out}\mathbf{r_{o}}_{l}.$
This choice of architecture with an intermediate layer is done to reduce the number of parameters, resulting from the 3-d tensor that maps from the attention layer in the $F(\mathbf{W}_{\text{net}},\mathbf{r}_{l})$ to the attention weights $\mathbf{W}_{\text{net}}$ mapping from a vector to a matrix.
@@ III.3 Transformer Architectures
Transformers [1] are fully trainable neural networks that have achieve state-of-the-art performance on many sequence modeling tasks. The input sequence of characters is mapped to embeddings of dimension $d$. A stack of $L$ layers follows, each containing a multi-head self-attention block and a feed-forward sub-block. Attention heads learn to focus on relevant parts of the input sequence at different positions. The feed-forward layers, typically parameterized by matrices of dimension on the order of $d\times\alpha d$ (where $\alpha$ is a constant expansion factor), refine the representation. The complexity of the transformer grows with $L$ (the number of layers) and $d$ (the embedding dimension). The output is passed through a final projection to logits of size $\mathcal{V}$, the vocabulary dimension. All parameters (including embedding, attention, feed-forward, and output projection) are learned end-to-end, giving transformers a large capacity to capture long-range dependencies. However, this capacity often comes at a significant computational and memory cost, particularly for large $L$, $d$ and long sequences.
IV Architectures
@@
The number of trainable parameters depends strongly on the chosen architecture and hyperparameters. We focused on choosing hyperparameters, such that the number of trainable parameters is comparable between all 3 different models. Table 1 summarizes the number of trainable parameters for each architecture across different configurations. With a vocabulary size of $V=59$ characters due to the Shakespeare dataset used in our experiments we arrive at architectures that range from around 15000 trainable parameters to around 150000.
For all three models, we take an embedding dimension of $d=16$. For the Transformer model we vary the hidden size $d_{h}$, the number of heads $h$, and the number of layers $L$. For the classic reservoir and AERC configurations, we vary the reservoir size $N$, the hidden dimensions of the attention layer $H$, and set the final layer $H_{o}=H$ to be equal to the attention layers hidden dimension.
to be, or not clear born have the poor my turn with my soul. romeo what when queen let this shall let the sbatter did the death, well the part, that your grief is turned, we hard me, like on an i when a true, and her-promain the commone to with you this lodged prithee to be, or not as a find. hele to where i wheress thieve hot may come, and you the sease, thy agains for your gon. pring thy worney that light and he winder, eyes was from three like io be good reator, but have them beenes of be strengler. i have such a foglingers widdly, good time i, btund. to be, or not my could now richment. lodovico-ling, as yield make one in their if younger, farewell. duchall, then, my great, teater these is and that beauty brow man, for here is the lord of play birthing on the had your and, but i’ll gench i heart that this mans colors well TransformerReservoirAERC
Figure 2: Comparison of language models: Transformer, Reservoir, and AERC with around 155k parameters (biggest model in Table 1) run in auto-regressive mode with the initial seed ”to be, or not” showing the generated text.
V Training
@@
where $\mathcal{V}$ is the number of output classes, $\hat{y}_{i}$ are the output logits, and $y_{i}$ is the true one-hot encoded label. We apply the softmax function to obtain output probabilities and compute the cross-entropy loss accordingly, a standard formulation in classification tasks [23]. All trainable weights are updated using backpropagation [24] and optimized using the Adam optimizer, a widely adopted variant of stochastic gradient descent [25].
VI Methodology
We divide the Shakespeare corpus into 666 shards. Five shards serve as training data, and one shard is used for testing. We sample sequences of length 323232, i.e. 32 characters as input, and train on a batch size of 102410241024 with Adam at a learning rate $1\times 10^{-4}$. The traditional reservoir, the attention-enhanced reservoir, and the transformer models are trained entirely by backpropagation in a consistent framework. We measure cross-entropy on the test shard as our main performance metric. We also assess closed-loop generation by feeding the predicted character back as input, measuring distributional metrics and repetitive patterns. Our parameter sweep includes varying the reservoir size or the neural network size in the attention reservoir, as well as varying the hidden dimension $d$ and number of layers and heads $L$ in the transformer. To save compute power, we first compute all the reservoir responses in one shard and keep them in memory, after which a few epochs on this shard are run before the next shard is processed. This procedure is repeated multiple times for all shards.
VII Results
We now present the performance of the trained models and analyze the training process. Figure 2 shows a representative example of text generated by each of the three models, all configured with approximately 155,000 trainable parameters—corresponding to the largest models listed in Table 1. While the generated text mimics the style of the Shakespearean corpus, the limited model capacity and small vocabulary size lead to inconsistencies and occasional errors. Nonetheless, the examples demonstrate that all three models successfully learn character-level prediction distributions that capture essential patterns in the dataset.
To quantitatively assess model performance, we report training and testing losses across epochs and data shards for all five model sizes described in Table 1. The results are visualized in Figure 3.

### v3

@@ Abstract
Large Language Models (LLM) have dominated the science and media landscape duo to their impressive performance on processing large chunks of data and produce human-like levels of text. Nevertheless, their huge energy demand and slow processing are still a bottleneck to further increasing quality while also making the models accessible to everyone. To solve this bottleneck, we will investigate how reservoir computing performs on natural text processing, which could enable fast and energy efficient hardware implementations. Studies investigating the use of reservoir computing as a language model remain sparse. In this paper, we compare three distinct approaches for character-level language modeling, two different reservoir computing approaches, where only an output layer is trainable, and the well-known transformer-based architectures, which fully learn an attention-based sequence representation. We explore the performance, computational cost and prediction accuracy for both paradigms by equally varying the number of trainable parameters for all models. Using a consistent pipeline for all three approaches, we demonstrate that transformers excel in prediction quality, whereas reservoir computers remain highly efficient reducing the training and inference speed. Furthermore, we investigate two types of reservoir computing: a traditional reservoir with a static linear readout, and an attention-enhanced reservoir that dynamically adapts its output weights via an attention mechanism. Our findings underline how these paradigms scale and offer guidelines to balance resource constraints with performance.
@@ I Introduction
Modern sequence modeling tasks, such as language modeling and machine translation, have been dominated by attention-based architectures—most prominently the transformer family [29, 5, 22, 2, 25, 7]—which excel at capturing long-range dependencies by learning contextual representations through self-attention layers and feed-forward networks. However, transformer training and inference incur substantial computational and energy costs, often requiring specialized hardware and limiting accessibility under tighter resource budgets.
@@
Figure 1: Diagram illustrating all three ML agent models with vector embedding, processing (Reservoir, AERC, Transformer), output layer, and predicted probability vector for the next letter. The vector embedding is trained for the transformer case, while being randomly initialized for the reservoirs. The predicted probabilities (e.g., $p_{A}$ for letter a) are used to compute the cross entropy loss $H(y,\hat{y})=-\sum_{i}y_{i}\log\hat{y}_{i}$, which drives the error feedback.
Reservoir computing (RC) offers an alternative computational paradigm. In RC, a large, fixed recurrent “reservoir” projects inputs into a high-dimensional state space, and only a lightweight readout layer is trained. This design dramatically reduces training time and energy consumption and can be implemented efficiently on analog hardware substrates. It has also been shown to excel in time-series prediction [9, 21, 16], capitalizing on its large, fixed random recurrent structures for capturing nonlinear dynamics with minimal training overhead [10, 17, 18, 26]. A key advantage of reservoir computing lies in its versatile physical realizations, spanning electronic, optical, and even quantum hardware [6, 8, 19, 4].
Recently, more expressive control over reservoir behavior has been introduced via neural programming paradigms, further expanding their utility in symbolic computation and structured sequence modeling [13], while other approaches introduced attention mechanisms at the readout stage to adaptively weight reservoir states [15]. While prior work has explored the use of reservoirs for text classification tasks [24] and even demonstrated hardware-oriented implementations for language learning [27], the use of reservoir computing explicitly as a language model—that is, for generative next-token prediction—remains largely unexplored.
Although transformers remain the dominant and most effective architecture for large-scale Natural Language Processing (NLP) [2], our goal is not to challenge this paradigm but to explore a complementary direction. Reservoir computing offers fundamentally different computational characteristics: fixed high-dimensional recurrent dynamics, strong energy efficiency, and inherent suitability for analog, neuromorphic, or photonic substrates. In settings where computational or energy budgets are constrained or full training of the underlying hardware is not feasible, such as embedded or edge devices, these properties make RC an appealing alternative. To reflect this motivation more clearly, the present work is framed as a proof-of-concept study investigating whether reservoir-based architectures can support next-token language modeling.
In this study, we present a unified framework, applying three distinct approaches — namely the classic RC, the attention-enhanced RC, and the transformer approach — to character-level sequence prediction on a small corpus of Shakespeare text inspired by recent minimalist transformer implementations such as Karpathy’s NanoGPT [12]. We vary reservoir size in the classic RC and RC size and attention layer size for an attention-enhanced readout [15], while varying hidden dimension and number of layers in transformers, obtaining multiple configurations spanning from thousands up to hundred thousand of trainable parameters. By measuring cross-entropy loss across these diverse setups, we expose fundamental trade-offs in resource usage and generalization performance.
In this work we additionally introduce a layered reservoir language model that adapts reservoir computing to modern token-level causal language modeling. By stacking multiple fixed recurrent reservoirs together with lightweight trainable mixing and feed-forward components, the model retains the depth and normalization structure characteristic of transformer architectures while replacing self-attention layers with computationally efficient recurrent dynamics. We evaluate this architecture at various model scales, compare it against transformer baselines under matched parameter budgets, and analyze both optimization behavior and scaling characteristics.
II Methods and Model Architectures
II.1 Traditional Reservoir Computing
@@
The echo state implementation of a classic reservoir computer consists of three main components: an input-to-reservoir weight matrix $\mathbf{W}_{\mathrm{in}}\in\mathbb{R}^{N\times d}$, a fixed recurrent reservoir matrix $\mathbf{W}_{\mathrm{res}}\in\mathbb{R}^{N\times N}$, and a readout layer $\mathbf{W}_{\mathrm{out}}\in\mathbb{R}^{\mathcal{V}\times N}$ [10]. Here $d$ is the dimension of the chosen input-embedding vector for the input letter, $N$ the number of nodes in the reservoir, and $\mathcal{V}$ the size of the vocabulary. We treat all state and input vectors as column vectors. The reservoir evolves its hidden states $\mathbf{r}_{t}\in\mathbb{R}^{N}$ according to
$\mathbf{r}_{t}=\tanh\!\big(\mathbf{W}_{\mathrm{res}}\,\mathbf{r}_{t-1}+\mathbf{W}_{\mathrm{in}}\,\mathbf{x}_{t}\big),$ (1)
@@
where $\mathbf{x}_{t}\in\mathbb{R}^{d}$ is the embedding vector of the current input character. In the simplest (traditional) reservoir setting, only the readout layer $\mathbf{W}_{\mathrm{out}}$ is trained:
$\bar{\mathbf{y}}_{t}=\mathbf{W}_{\mathrm{out}}\,\mathbf{r}_{t},$ (2)
@@
where $\bar{\mathbf{y}}_{t}\in\mathbb{R}^{\mathcal{V}}$ denotes the unnormalized logits at time $t$. Typically, $\mathbf{W}_{\mathrm{out}}$ is learned using ridge regression, which minimizes the mean squared error with an added $\ell_{2}$ regularization term to prevent overfitting and improve generalization. However, in our simulations, we employ a cross-entropy loss function, necessitating gradient-based optimization methods for training $\mathbf{W}_{\mathrm{out}}$. Moreover, ridge regression can become computationally intensive as the number of training samples and reservoir size increase, due to the inversion of large matrices. By using gradient-based methods, we circumvent these scalability issues.
@@
Table 1: Trainable Parameter Counts for the classic reservoir, AERC, and Transformer Architectures. All systems used an embedding dimension of 16 for the 59 Vocabularies.
@@
Transformer Hidden Size $d_{h}=64$, Heads $h=4$, Layers $L=4$ $15067$
Hidden Size $d_{h}=72$, Heads $h=8$, Layers $L=8$ $30299$
Hidden Size $d_{h}=128$, Heads $h=8$, Layers $L=8$ $45083$
Hidden Size $d_{h}=356$, Heads $h=8$, Layers $L=8$ $105275$
Hidden Size $d_{h}=256$, Heads $h=16$, Layers $L=16$ $155803$
Reservoir Reservoir Size $N=250$ $14809$
Reservoir Size $N=500$ $29559$
Reservoir Size $N=750$ $44309$
Reservoir Size $N=1750$ $103309$
Reservoir Size $N=2600$ $153459$
Att Reservoir Reservoir Size $N=75$, Hidden Size $H=13$ $15464$
Reservoir Size $N=75$, Hidden Size $H=19$ $31124$
Reservoir Size $N=100$, Hidden Size $H=20$ $45259$
Reservoir Size $N=150$, Hidden Size $H=25$ $102809$
Reservoir Size $N=160$, Hidden Size $H=30$ $155459$
II.2 Attention-Enhanced Reservoir Computing
@@
$\mathbf{W}_{\text{att},l}=F(\mathbf{W}_{\text{net}},\mathbf{r}_{l}).$ (3)
@@
$\mathbf{r_{o}}_{l}=\mathbf{W}_{\text{att},l}\mathbf{r}_{l}.$ (4)
@@
$\bar{\mathbf{y}}_{l}=\mathbf{W}_{out}\mathbf{r_{o}}_{l}.$ (5)
This choice of architecture with an intermediate layer is done to reduce the number of parameters.
@@ II.3 Transformer Architectures
Transformers [29] are fully trainable neural networks that have achieve state-of-the-art performance on many sequence modeling tasks. The input sequence of characters is mapped to embeddings of dimension $d$. A stack of $L$ layers follows, each containing a multi-head self-attention block and a feed-forward sub-block. Attention heads learn to focus on relevant parts of the input sequence at different positions. The feed-forward layers, typically parameterized by matrices of dimension on the order of $d_{h}=d\times\alpha d$ (where $\alpha$ is a constant expansion factor), refine the representation. The complexity of the transformer grows with $L$ (the number of layers) and $d$ (the embedding dimension). The output is passed through a final projection to logits of size $\mathcal{V}$, the vocabulary dimension. All parameters (including embedding, attention, feed-forward, and output projection) are learned end-to-end, giving transformers a large capacity to capture long-range dependencies. However, this capacity often comes at a significant computational and memory cost, particularly for large $L$, $d$ and long sequences.
In standard transformer architectures, the expansion factor $\alpha$ is typically chosen between $2$ and $4$, making the intermediate feed-forward dimension $d_{h}=\alpha d$ larger than the embedding dimension. In our simulations, however, we keep the embedding dimension fixed at $d=16$ and vary only the internal feed-forward dimensions to ensure that all architectures—Transformer, classic reservoir, and AERC—operate under comparable embedding dimensions. This design choice enables a fair comparison of representational efficiency across models without conflating performance differences with changes in embedding dimensionality.
II.4 Architectures
@@
The number of trainable parameters depends strongly on the chosen architecture and hyperparameters. We focused on choosing hyperparameters, such that the number of trainable parameters is comparable between all 3 different models. Table 1 summarizes the number of trainable parameters for each architecture across different configurations. With a vocabulary size of $V=59$ characters due to the Shakespeare dataset used in our simulations we arrive at architectures that range from around 15000 trainable parameters to around 150000.
For all three models, we take an embedding dimension of $d=16$. For the Transformer model we vary the hidden size $d_{h}$, the number of heads $h$, and the number of layers $L$. For the classic reservoir and AERC configurations, we vary the reservoir size $N$, the hidden dimensions of the attention layer $H$, and set the final layer $H_{o}=H$ to be equal to the attention layers hidden dimension. In the reservoir-based architectures (classic reservoir and AERC), only the readout and gating networks are trained using gradient descent. The internal reservoir matrices $W_{\mathrm{in}}$ and $W_{\mathrm{res}}$ are initialized once and remain fixed throughout training; they are therefore not counted as trainable parameters in Table 1. This definition is consistent with the long-term perspective of reservoir computing, where the reservoir dynamics are viewed as a fixed free computation implemented directly in analog or photonic hardware—while only digital components are optimized.
II.5 Training
@@
where $\mathcal{V}$ is the number of output classes, $\hat{y}_{i}$ are the output probabilities, and $y_{i}$ is the true one-hot encoded label. We apply the softmax function to obtain output probabilities and compute the cross-entropy loss accordingly, a standard formulation in classification tasks [1]. All trainable weights are updated using backpropagation [23] and optimized using the Adam optimizer, a widely adopted variant of stochastic gradient descent [14].
II.6 Dataset and Methodology
We study a character-level next-character prediction task. Let $\mathbf{X}=\{x_{1},\dots,x_{T}\}$ denote a fixed-length input sequence of characters, and let $y_{T+1}$ be the next character to predict. The task is formulated as a multi-class classification problem over a vocabulary of size $\mathcal{V}=59$.
We use a 9-million-character corpus of Shakespeare text, inspired by minimalist transformer implementations such as Karpathy’s NanoGPT [12]. All text is lowercased during preprocessing to reduce the vocabulary size. The corpus is split into six contiguous shards: five shards for training and one shard for testing. Each shard is processed at the character level, yielding sequences of length $32$ as input, with the subsequent character serving as the prediction target.
Training minimizes the standard cross-entropy loss between predicted logits and the ground-truth next character. Models are trained using Adam with a learning rate of $1\times 10^{-4}$ and a batch size of $1024$. We track both training and test losses per shard and per epoch, reporting cross-entropy on the held-out test shard as the primary performance metric.
All models—the traditional reservoir, the attention-enhanced reservoir, and the transformer baselines—are trained end-to-end via backpropagation within a unified training framework. In addition to next-character prediction accuracy, we evaluate closed-loop text generation by feeding predicted characters back as input and analyzing distributional statistics and repetitive patterns. Our parameter sweep varies the reservoir size or neural subnetwork size in the attention-based reservoir models, as well as the hidden dimension $d$ and number of layers $L$ (and heads) in the transformer models.
To reduce computational cost, reservoir responses are precomputed shard-wise and kept in memory, allowing several training epochs to be run on each shard before advancing to the next. This process is repeated cyclically across all shards.
All simulations were conducted on a single NVIDIA RTX 4090 GPU running Ubuntu 22.04, using PyTorch 2.4.1 with CUDA 12.1. Mixed-precision training (bf16/fp16 with loss scaling) was used throughout. The full codebase, configuration files, and simulation scripts are publicly available at (See Data Availability).
III Results
III.1 Comparison of Training and Testing Loss
We now present the performance of the trained models and analyze the training process.